OpenStudy (anonymous):

Evaluate the indefinite integral: xsin(x^2)dx We are supposed to use U substitution and i substituted everything in but I'm not sure I substituted du correctly.

OpenStudy (anonymous):

i would try using u = x^2 .

OpenStudy (anonymous):

i did and then i made du=xsinsdx but when I put it all together it is not making sense

OpenStudy (anonymous):

-0.5cos(x^2)

OpenStudy (anonymous):

let u = x^2. du = 2x dx or x dx = du /2. then the integration will be sin (u) du/2

OpenStudy (anonymous):

thank you both!

OpenStudy (anonymous):

welcome ;)